A normally distributed random variable X has a mean of 500 and a standard deviation of 40.


1
Expert's answer
2021-03-16T04:54:16-0400

Q: P(X>530)

Mean("\mu" )= 500

Standard deviation("\sigma" )= 40


Normal distribution formula"\Rightarrow z=\dfrac{\bar x-\mu}{\sigma}"


"z=\dfrac{530-500}{40}= \dfrac{30}{40}=0.75"

Now, P(z>0.75)=0.5-0.2734=0.2266

Hence, P(X>530)=0.2266